# An improvement toward global boundedness in a fully parabolic chemotaxis with singular sensitivity in any dimension

### v1

@@ 3 A priori estimates
In this section, we establish several estimates for solutions to (1.1), which will be later applied to derive the most essential estimate, Lemma 4.1, allowing to obtain an $L^{\frac{n}{2}+}$ bound for $u$. Let us commence with the following lemma enabling us to control the quantity $\int_{\Omega}v^{p-r}$.
@@
For any $p>1$, $10<r<p-$ 1, $M>0$ and $\epsilon>0$, there exists $C=C(p,r,M,\epsilon,u_{0},v_{0})>0$ such that
$\frac{d}{dt}\int_{\Omega}v^{p-r}+M\int_{\Omega}v^{p-r}\leq\epsilon\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}+C,$
for all $t\in(0,T_{\rm max})$.
@@
Applying Young’s inequality with $\epsilon>0$ deduces that
@@
$\leq\epsilon\int_{\Omega}\frac{u^{p+1}}{v^{p-r}}+c_{1}\int_{\Omega}v^{p-r},$
where $c_{1}=\epsilon^{-\frac{1}{p}}(p-r)^{\frac{p+1}{p}}$. Thanks to Gagliardo–Nirenberg interpolation inequality, Young’s inequality, and Lemma 2.2 we obtain that
$\left(M+c_{1}-\frac{p+1}{2}\right)_{+}\int_{\Omega}v^{p-r}$ $\leq c_{2}\left(\int_{\Omega}|\nabla v^{\frac{p-r}{2}}|^{2}\right)^{\theta}\left(\int_{\Omega}v\right)^{(p-r)(1-\theta)}+c_{2}\left(\int_{\Omega}v\right)^{p-r}$
@@
where $\theta=\frac{\frac{p-r}{2}-\frac{1}{2}}{\frac{p-r}{2}+\frac{1}{n}-\frac{1}{2}}\in(0,1)$, $c_{2}>0$, and $c_{3}>0$. Finally, combining (3.1) and (3) proves the desired result. ∎
@@
Let $p>1$ and $r=\frac{p-1}{2}$. Then there exists $C=C(p,n)>0$ such that the following holds
@@
$\quad-\frac{p-1}{4}\int_{\Omega}v^{-p-r-2}|\nabla v|^{2p+2}+C\int_{\Omega}v^{p-r}$
$\quad+2p(2p-2+\sqrt{n})\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-2}|D^{2}v|$
$\quad+(2p-1)(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u$ (3.3)
@@
By using of the second equation of (1.1) as well as the identities $\nabla v\cdot\nabla\Delta v=\frac{1}{2}\Delta|\nabla v|^{2}-|D^{2}v|^{2}$ and $\nabla|\nabla v|^{2}=2D^{2}v\cdot\nabla v$, and several integration by parts, we obtain that
@@
$\quad+2p(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u-2p(p-1)\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-4}\nabla|\nabla v|^{2}\cdot\nabla v$
$\quad-2p\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-2}\Delta v$
@@
$\quad+p\int_{\partial\Omega}v^{-p-r}|\nabla v|^{2p-2}\cdot\frac{\partial|\nabla v|^{2}}{\partial\nu}-2p\int_{\Omega}v^{-p-r}|\nabla v|^{2p-2}|D^{2}v|^{2}-2p\int_{\Omega}v^{-p-r}|\nabla v|^{2p}$
$\quad+2p(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u-4p(p-1)\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-4}\nabla v\cdot D^{2}v\cdot\nabla v$
$\quad-2p\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-2}\Delta v,$ (3.5)
@@
$\quad+(p+r)\int_{\Omega}v^{-p-r}|\nabla v|^{2p}-(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u.$ (3.6)
@@
$\quad+2p(p+r-1)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p-2}\nabla|\nabla v|^{2}\cdot\nabla v-(p-r)\int_{\Omega}v^{-p-r}|\nabla v|^{2p}$
$\quad-\left[(p+r)(p+r+1)-2p\right]\int_{\Omega}v^{-p-r-2}|\nabla v|^{2p+2}+(2p-1)(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u$
$\quad-4p(p-1)\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-4}\nabla v\cdot D^{2}v\cdot\nabla v-2p\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-2}\Delta v$
$\quad+p\int_{\partial\Omega}v^{-p-r}|\nabla v|^{2p-2}\cdot\frac{\partial|\nabla v|^{2}}{\partial\nu}.$
Applying the inequality that $|\Delta v|\leq\sqrt{n}|D^{2}v|$ and Young’s inequality yields that
@@
$\quad-\left[(p+r)(p+r+1)-2p-\frac{p(p+r-1)^{2}}{p-1}\right]\int_{\Omega}v^{-p-r-2}|\nabla v|^{2p+2}$
$\quad+2p(2p-2+\sqrt{n})\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-2}|D^{2}v|$
$\quad+(2p-1)(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u+p\int_{\partial\Omega}v^{-p-r}|\nabla v|^{2p-2}\cdot\frac{\partial|\nabla v|^{2}}{\partial\nu}.$ (3.7)
Using Lemma 2.6 and noting that $(p+r)(p+r+1)-2p-\frac{p(p+r-1)^{2}}{p-1}=\frac{p-1}{4}$ since $r=\frac{p-1}{2}$, (3) deduces
@@
$\quad-\frac{p-1}{4}\int_{\Omega}v^{-p-r-2}|\nabla v|^{2p+2}+c_{1}\int_{\Omega}v^{p-r}$
$\quad+2p(2p-2+\sqrt{n})\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-2}|D^{2}v|$
$\quad+(2p-1)(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u,$ (3.8)
@@
Let $p\in\left(\frac{n}{2},n\right)$ and $r=\frac{p-1}{2}$. Then there exist positive constants $C=C(p,n)>0$ and $L_{1}$, $L_{2}$ and $L_{3}$ depending only on $n$ such that
@@
for all $t\in(0,T_{\rm max})$.
@@
Using Lemma 3.2 and noting that $\frac{n}{2}<p<n$ yields that
@@
where the last inequality holds due to
$\frac{p+r+1+\sqrt{n}}{p-r+1}=\frac{3p+1+2\sqrt{n}}{p+3}<5.$ (3.12)
Similarly, we also have
@@
where the last inequality is true because of (3.12) and
$\frac{p-r+1}{p-r-1}=\frac{p+3}{p-1}<9.$
In light of Young’s inequality, we obtain that
$c_{3}\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u$ $\leq\frac{n-2}{16}\int_{\Omega}\frac{|\nabla v|^{2p+2}}{v^{p+r+2}}+\left(\frac{16}{n-2}\right)^{p}c_{3}^{p+1}\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}$
@@
Applying Hölder’s inequality, (3), (3) and Young’s inequality entails that
@@
Finally, combining (3), (3) and (3) deduces (3.9) with $L_{1}=\frac{n+2}{8}$, $L_{2}=\frac{n-2}{16}$ and $L_{3}=16^{n}c_{3}^{n+1}+(152c_{2})^{n+1}$. ∎
@@
Let $p\in\mathbb{R}$ and $r\in\mathbb{R}$. Then the following identity hold for any $t\in(0,T_{\rm max})$:
@@
$\quad+\left[2pr+p(p-1)\chi\right]\int_{\Omega}u^{p-1}v^{-r-1}\nabla u\cdot\nabla v+r\int_{\Omega}u^{p}v^{-r}.$
@@
There exist positive constants $L_{4}$ and $L_{5}$ depending only on $n$ such that for any $p\in\left(\frac{n}{2},n\right)$ the following holds
$\frac{d}{dt}\int_{\Omega}u^{p}v^{-r}+\int_{\Omega}u^{p}v^{-r}+L_{4}\int_{\Omega}u^{p+1}v^{-r-1}\leq L_{5}(\chi^{2}p-1)_{+}\int_{\Omega}u^{p}v^{-p-r-2}|\nabla v|^{2}+C\int_{\Omega}v^{p-r},$ (3.16)
where $r=\frac{p-1}{2}$ and $C=C(p,n)>0$.
@@
$\quad+\left[2pr+p(p-1)\chi\right]\int_{\Omega}u^{p-1}v^{-r-1}\nabla u\cdot\nabla v+r\int_{\Omega}u^{p}v^{-r}.$ (3.17)
@@
$\quad+\frac{\left[2pr+p(p-1)\chi\right]^{2}}{4p(p-1)}\int_{\Omega}u^{p}v^{-r-2}|\nabla v|^{2}.$
@@
Collecting (3), (3.18) and (3.19) and noting that $\frac{r}{2}>\frac{n-2}{8}$ deduces (3.16) with $L_{4}=\frac{n-2}{8}$ and $L_{5}=\frac{n-1}{4}$, which completes the proof.
@@ 4 Proof of the main result
Let us begin this section by showing that the energy function $F_{\lambda}(u,v)$ with an appropriate choice for $\lambda$ is uniformly bounded.
@@
For any $n\geq 3$, there exists $\delta=\delta(n)>0$ such that for any $0<\chi<\sqrt{\frac{2(1+\delta)}{n}}$ and $1<p<\min\left\{\frac{1+\delta}{\chi^{2}},n\right\}$, there exists a constant $C=C(u_{0},v_{0},n,\Omega,p)>0$ such that
$\int_{\Omega}v^{p-r}+\int_{\Omega}u^{p}v^{-r}+\int_{\Omega}v^{-p-r}|\nabla v|^{2p}\leq C,\qquad\text{for all }t\in(0,T_{\rm max}),$ (4.1)
where $r=\frac{p-1}{2}$.
@@
We assume first that $p>\frac{n}{2}$, and define $y(t):=\int_{\Omega}u^{p}v^{-r}+\lambda\int_{\Omega}v^{-p-r}|\nabla v|^{2p}$, where $\lambda=\frac{L_{4}}{2L_{3}}$. From Lemma 3.3 and Lemma 3.5, we find that
@@
$\quad+c_{2}\int_{\Omega}v^{p-r}.$ (4.2)
$c_{1}=\min\left\{1,\frac{L_{1}L_{4}}{2L_{3}}\right\}$ and $c_{2}>0$. Applying Young’s inequality entails that
@@
$\leq L_{5}\left(\chi^{2}p-1\right)_{+}\int_{\Omega}u^{p+1}v^{-r-1}+L_{5}\left(\chi^{2}p-1\right)_{+}\int_{\Omega}\frac{|\nabla v|^{2p+2}}{v^{p+r+2}}.$ (4.3)
Denoting $\delta=\min\left\{\frac{L_{2}L_{4}}{2L_{3}L_{5}},\frac{L_{4}}{2L_{5}}\right\}>\frac{(n-2)^{2}}{256(n-1)\left(16^{n}(3n^{2})^{n+1}+(912n^{2})^{n+1}\right)}$, then for any $\frac{n}{2}<p<\frac{1+\delta}{\chi^{2}}$ we have that $L_{5}\left(\chi^{2}p-1\right)_{+}\leq\min\left\{\frac{L_{2}L_{4}}{2L_{3}},\frac{L_{4}}{4}\right\}$. This, together with (4) and (4) implies that
$y^{\prime}(t)+c_{1}y(t)+\frac{L_{4}}{4}\int_{\Omega}u^{p+1}v^{-r-1}\leq c_{2}\int_{\Omega}v^{p-r},\qquad\text{for all }t\in(0,T_{\rm max}).$ (4.4)
Applying Lemma 3.1 with $M=c_{2}+c_{1}$ and $\epsilon=\frac{L_{4}}{4}$ entails that
$\frac{d}{dt}\int_{\Omega}v^{p-r}+(c_{2}+c_{1})\int_{\Omega}v^{p-r}\leq\frac{L_{4}}{4}\int_{\Omega}u^{p+1}v^{-r-1}+c_{3},$ (4.5)
for some $c_{3}>0$. Combining (4.4) and (4.5) implies that
$\frac{d}{dt}F_{\lambda}(u,v)+c_{1}F_{\lambda}(u,v)\leq c_{3}\qquad\text{for all }t\in(0,T_{\rm max}),$
which combines with Gronwall’s inequality deduces (4.1) for any $\frac{n}{2}<p<\frac{1+\delta}{\chi^{2}}$. Now, for any $1<q<p$, applying Hölder’s inequality and Lemma 2.2 entails that
@@
Applying Lemma 4.1 yields that there exists $p_{0}\in\left(\frac{n}{2},n\right)$ such that
@@
then we see from Lemma 2.1 that $M(T)<\infty$. Since $p_{0}>\frac{n}{2}$ implies that $\frac{p_{0}+1}{2(n-p_{0}+1)}>\frac{n}{2}$, we can choose $p\in\left(\frac{n}{2},\min\left\{p_{0},\frac{p_{0}+1}{2(n-p_{0}+1)}\right\}\right)$. Employing Hölder’s inequality entails that
$\int_{\Omega}u^{p}\leq\left(\int_{\Omega}u^{p_{0}}v^{-r}\right)^{\frac{p}{p_{0}}}\cdot\left(\int_{\Omega}v^{\frac{pr}{p_{0}-p}}\right)^{\frac{p_{0}-p}{p_{0}}},$
which combines with (4.6) implies that
$\left\|u(\cdot,t)\right\|_{L^{p}(\Omega)}\leq c_{1}^{\frac{1}{p_{0}}}\left\|v(\cdot,t)\right\|_{L^{\frac{pr}{p_{0}-p}}(\Omega)}^{\frac{r}{p_{0}}}\qquad\text{for all }t\in(0,T_{\rm max}).$ (4.7)

### v2

@@ 3 A priori estimates
In this section, we establish several estimates for solutions to (1.1), which will be later applied to derive the most essential estimate, Lemma 4.1, allowing us to obtain a uniform in time $L^{p}$ bound with some $p>\frac{n}{2}$ for $u$. Let us commence with the following lemma, which enables us to control the quantity $\int_{\Omega}v^{p-r}$.
@@
For any $p>1$, 0<r<p10<r<p-1, $M>0$, $\chi>0$ and $\epsilon>0$, there exists $C=C(p,r,M,\epsilon,u_{0},v_{0})>0$ such that
$\frac{d}{dt}\int_{\Omega}v^{p-r}+M\int_{\Omega}v^{p-r}\leq\epsilon\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}+C\hskip 18.49988pt\text{for all }t\in(0,T_{\rm max}).$
@@
Applying Young’s inequality with $\epsilon>0$ yields
@@
$\leq\epsilon\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}+c_{1}\int_{\Omega}v^{p-r},$
where $c_{1}=\epsilon^{-\frac{1}{p}}(p-r)^{\frac{p+1}{p}}$. This, together with (3.1) implies that
$\frac{d}{dt}\int_{\Omega}v^{p-r}+M\int_{\Omega}v^{p-r}$ $\leq-(p-r)(p-r-1)\int_{\Omega}v^{p-r-2}|\nabla v|^{2}+\epsilon\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}$
$\qquad+(M+c_{1}-{p+r})\int_{\Omega}v^{p-r}\hskip 18.49988pt\text{for all }t\in(0,T_{\rm max}).$ (3.2)
Thanks to Gagliardo–Nirenberg interpolation inequality, Young’s inequality, and Lemma 2.2, it follows that
$\left(M+c_{1}-{p+r}\right)_{+}\int_{\Omega}v^{p-r}$ $\leq c_{2}\left(\int_{\Omega}|\nabla v^{\frac{p-r}{2}}|^{2}\right)^{\theta}\left(\int_{\Omega}v\right)^{(p-r)(1-\theta)}+c_{2}\left(\int_{\Omega}v\right)^{p-r}$
@@
where $\theta=\frac{\frac{p-r}{2}-\frac{1}{2}}{\frac{p-r}{2}+\frac{1}{n}-\frac{1}{2}}\in(0,1)$, $c_{2}>0$, and $c_{3}>0$. Finally, combining (3) and (3), we arrive at
$\frac{d}{dt}\int_{\Omega}v^{p-r}+M\int_{\Omega}v^{p-r}\leq\epsilon\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}+c_{3}\hskip 18.49988pt\text{for all }t\in(0,T_{\rm max}),$
which completes the proof.
@@
Let $p>1$, $r=\frac{p-1}{2}$ and $\chi>0$. Then there exists $C=C(p,n)>0$ such that the following holds
@@
$\qquad-\frac{p-1}{4}\int_{\Omega}v^{-p-r-2}|\nabla v|^{2p+2}+C\int_{\Omega}v^{p-r}$
$\qquad+2p(2p-2+\sqrt{n})\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-2}|D^{2}v|$
$\qquad+(2p-1)(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u$ (3.4)
@@
By using the second equation of (1.1) as well as the identities $\nabla v\cdot\nabla\Delta v=\frac{1}{2}\Delta|\nabla v|^{2}-|D^{2}v|^{2}$ and $\nabla|\nabla v|^{2}=2D^{2}v\cdot\nabla v$, and performing several integration by parts, we obtain that
@@
$\qquad+2p(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u-2p(p-1)\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-4}\nabla|\nabla v|^{2}\cdot\nabla v$
$\qquad-2p\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-2}\Delta v$
@@
$\qquad+p\int_{\partial\Omega}v^{-p-r}|\nabla v|^{2p-2}\cdot\frac{\partial|\nabla v|^{2}}{\partial\nu}-2p\int_{\Omega}v^{-p-r}|\nabla v|^{2p-2}|D^{2}v|^{2}-2p\int_{\Omega}v^{-p-r}|\nabla v|^{2p}$
$\qquad+2p(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u-4p(p-1)\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-4}\nabla v\cdot D^{2}v\cdot\nabla v$
$\qquad-2p\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-2}\Delta v,$ (3.6)
@@
$\qquad+(p+r)\int_{\Omega}v^{-p-r}|\nabla v|^{2p}-(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u.$ (3.7)
@@
$\qquad+2p(p+r-1)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p-2}\nabla|\nabla v|^{2}\cdot\nabla v-(p-r)\int_{\Omega}v^{-p-r}|\nabla v|^{2p}$
$\qquad-\left[(p+r)(p+r+1)-2p\right]\int_{\Omega}v^{-p-r-2}|\nabla v|^{2p+2}+(2p-1)(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u$
$\qquad-4p(p-1)\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-4}\nabla v\cdot D^{2}v\cdot\nabla v-2p\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-2}\Delta v$
$\qquad+p\int_{\partial\Omega}v^{-p-r}|\nabla v|^{2p-2}\cdot\frac{\partial|\nabla v|^{2}}{\partial\nu}.$
Applying the inequality that $|\Delta v|\leq\sqrt{n}|D^{2}v|$ and Young’s inequality yields
@@
$\qquad-\left[(p+r)(p+r+1)-2p-\frac{p(p+r-1)^{2}}{p-1}\right]\int_{\Omega}v^{-p-r-2}|\nabla v|^{2p+2}$
$\qquad+2p(2p-2+\sqrt{n})\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-2}|D^{2}v|$
$\qquad+(2p-1)(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u+p\int_{\partial\Omega}v^{-p-r}|\nabla v|^{2p-2}\cdot\frac{\partial|\nabla v|^{2}}{\partial\nu}.$ (3.8)
Applying Lemma 2.6, noting that $(p+r)(p+r+1)-2p-\frac{p(p+r-1)^{2}}{p-1}=\frac{p-1}{4}$ since $r=\frac{p-1}{2}$, and using (3), we arrive at
@@
$\qquad-\frac{p-1}{4}\int_{\Omega}v^{-p-r-2}|\nabla v|^{2p+2}+c_{1}\int_{\Omega}v^{p-r}$
$\qquad+2p(2p-2+\sqrt{n})\int_{\Omega}uv^{-p-r}|\nabla v|^{2p-2}|D^{2}v|$
$\qquad+(2p-1)(p+r)\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u,$ (3.9)
@@
Let $p\in\bigl(\frac{n}{2},\,n\bigr)$, set $r=\frac{p-1}{2}$, and let $\chi>0$. Then there exists a constant $C=C(p,n)>0$ such that the inequality
@@
holds for all $t\in(0,T_{\max})$, where $L_{1}=\frac{n+2}{8}$ and the constants $L_{2}$ and $L_{3}$ are specified in Remark 1.1.
@@
Using Lemma 3.2 and noting that $\frac{n}{2}<p<n$ yields
@@
where the last inequality follows from
$\frac{p+r+1+\sqrt{n}}{p-r+1}=\frac{3p+1+2\sqrt{n}}{p+3}<5,$ (3.13)
which holds for all $p>\sqrt{n}-7$; this condition is satisfied since $p>\frac{n}{2}>\sqrt{n}-7$. Similarly, we also have
@@
where the last inequality follows from (3.13) together with the estimate
$\frac{p-r+1}{p-r-1}=\frac{p+3}{p-1}<9,$
which is valid since $p>\frac{n}{2}\geq\frac{3}{2}$. In light of Young’s inequality, we derive that
$c_{3}\int_{\Omega}v^{-p-r-1}|\nabla v|^{2p}u$ $={\int_{\Omega}\left[\left(\frac{n-2}{16}\right)^{\frac{p}{p+1}}\frac{|\nabla v|^{2p}}{v^{\frac{p(p+r+2)}{p+1}}}\right]\cdot\left[c_{3}\left(\frac{16}{n-2}\right)^{\frac{p}{p+1}}\frac{u}{v^{\frac{r+1}{p+1}}}\right]}$
${\leq\frac{p(n-2)}{16(p+1)}\int_{\Omega}\frac{|\nabla v|^{2p+2}}{v^{p+r+2}}+\left(\frac{16}{n-2}\right)^{p}\frac{c_{3}^{p+1}}{p+1}\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}}$
$\leq\frac{n-2}{16}\int_{\Omega}\frac{|\nabla v|^{2p+2}}{v^{p+r+2}}+\left(\frac{16}{n-2}\right)^{p}c_{3}^{p+1}\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}$
@@
Applying Hölder’s inequality, (3), (3) and Young’s inequality and noting that $p<n$ and $\sqrt{226}(102)^{\frac{p-1}{2(p+1)}}<152$, we obtain that
@@
${\leq\sqrt{226}(102)^{\frac{p-1}{2(p+1)}}c_{2}\left(\int_{\Omega}v^{-p-r+2}|\nabla v|^{2p-2}|D^{2}\ln v|^{2}\right)^{\frac{p}{p+1}}\left(\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}\right)^{\frac{1}{p+1}}}$
@@
${\leq\frac{np}{2(p+1)}\int_{\Omega}v^{-p-r+2}|\nabla v|^{2p-2}|D^{2}\ln v|^{2}+\frac{(152c_{2})^{p+1}}{p+1}\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}}$
@@
Combining (3), (3) and (3), we infer that
$\frac{d}{dt}\int_{\Omega}\frac{|\nabla v|^{2p}}{v^{p+r}}+\frac{n+2}{8}\int_{\Omega}\frac{|\nabla v|^{2p}}{v^{p+r}}+\frac{n-2}{16}\int_{\Omega}\frac{|\nabla v|^{2p+2}}{v^{p+r+2}}$ $\leq\left(16^{n}c_{3}^{n+1}+(152c_{2})^{n+1}\right)\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}$
$\qquad+c_{1}\int_{\Omega}v^{p-r}\hskip 18.49988pt\text{for all }t\in(0,T_{\rm max}).$
This implies (3.10) and thereby completes the proof. ∎
@@
Let $p\in\mathbb{R}$, $r\in\mathbb{R}$ and $\chi>0$. Then the following identity holds for any $t\in(0,T_{\rm max})$:
@@
$\qquad+\left[2pr+p(p-1)\chi\right]\int_{\Omega}u^{p-1}v^{-r-1}\nabla u\cdot\nabla v+r\int_{\Omega}u^{p}v^{-r}.$
@@
Let $r=\frac{p-1}{2}$ and $\chi>0$. There exist a positive constant $C=C(p,n)$ such that for any $p\in\left(\frac{n}{2},n\right)$ the following holds
$\frac{d}{dt}\int_{\Omega}u^{p}v^{-r}+\int_{\Omega}u^{p}v^{-r}+L_{4}\int_{\Omega}u^{p+1}v^{-r-1}$ $\leq L_{5}(\chi^{2}p-1)_{+}\int_{\Omega}u^{p}v^{-p-r-2}|\nabla v|^{2}$
$\qquad+C\int_{\Omega}v^{p-r}\hskip 18.49988pt\text{for all }t\in(0,T_{\rm max}),$ (3.17)
where $L_{4}$ and $L_{5}$ are given in Remark 1.1.
@@
$\qquad+\left[2pr+p(p-1)\chi\right]\int_{\Omega}u^{p-1}v^{-r-1}\nabla u\cdot\nabla v+r\int_{\Omega}u^{p}v^{-r}.$ (3.18)
@@
$\qquad+\frac{\left[2pr+p(p-1)\chi\right]^{2}}{4p(p-1)}\int_{\Omega}u^{p}v^{-r-2}|\nabla v|^{2}.$
@@
By combining (3)–(3.20) and noting that $\frac{r}{2}>\frac{n-2}{8}$ and $\frac{p-1}{4}>\frac{n-2}{8}$, we arrive at
$\frac{d}{dt}\int_{\Omega}u^{p}v^{-r}+\int_{\Omega}u^{p}v^{-r}+\frac{n-2}{8}\int_{\Omega}u^{p+1}v^{-r-1}$ $\leq\frac{(n-2)(\chi^{2}p-1)_{+}}{8}\int_{\Omega}u^{p}v^{-r-2}|\nabla v|^{2}$
$\qquad+c_{1}\int_{\Omega}v^{p-r}\hskip 18.49988pt\text{for all }t\in(0,T_{\rm max}).$
This proves (3.5) and thereby concludes the proof.
@@ 4 Proof of the main result
Let us begin this section by showing that the energy function $F_{\lambda}(u,v)$, with an appropriate choice of $\lambda$ is uniformly bounded.
@@
Let $r=\frac{p-1}{2}$ and $\delta=\frac{L_{2}L_{4}}{2L_{3}L_{5}},$ where $L_{i}$, with $i=2,3,4,5$ are given in Remark 1.1. For any
$0<\chi<\sqrt{\frac{2(1+\delta)}{n}}\quad\text{and}\quad 1<p<\min\left\{\frac{1+\delta}{\chi^{2}},\,n\right\},$
there exists a constant $C=C(u_{0},v_{0},n,\Omega,p)>0$ such that
$\int_{\Omega}v^{p-r}+\int_{\Omega}u^{p}v^{-r}+\int_{\Omega}v^{-p-r}|\nabla v|^{2p}\leq C\hskip 18.49988pt\text{for all }t\in(0,T_{\max}).$ (4.1)
@@
We first assume that $p>\frac{n}{2}$, and define $y(t):=\int_{\Omega}u^{p}v^{-r}+\lambda\int_{\Omega}v^{-p-r}|\nabla v|^{2p}$ with $\lambda=\frac{L_{4}}{2L_{3}}$. From Lemma 3.3 and Lemma 3.5, we find that
@@
$\qquad+c_{2}\int_{\Omega}v^{p-r}.$ (4.2)
$c_{1}=\min\left\{1,\frac{L_{1}L_{4}}{2L_{3}}\right\}$ with $L_{1}$ is given in Lemma 3.3 and $c_{2}>0$. Since $\delta=\frac{L_{2}L_{4}}{2L_{3}L_{5}}$, we infer that for any
$\frac{n}{2}<p<\frac{1+\delta}{\chi^{2}},$
it holds that
$L_{5}\big(\chi^{2}p-1\big)_{+}\leq\min\left\{\frac{L_{2}L_{4}}{2L_{3}},\,\frac{L_{4}}{4}\right\}.$
Using this and applying Young’s inequality yields
@@
$\leq{\frac{pL_{5}\left(\chi^{2}p-1\right)_{+}}{p+1}\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}+\frac{L_{5}\left(\chi^{2}p-1\right)_{+}}{p+1}\int_{\Omega}\frac{|\nabla v|^{2p+2}}{v^{p+r+2}}}$
$\leq L_{5}\left(\chi^{2}p-1\right)_{+}\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}+L_{5}\left(\chi^{2}p-1\right)_{+}\int_{\Omega}\frac{|\nabla v|^{2p+2}}{v^{p+r+2}}$
${\leq\frac{L_{4}}{4}\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}+\frac{L_{4}L_{2}}{2L_{3}}\int_{\Omega}\frac{|\nabla v|^{2p+2}}{v^{p+r+2}}.}$ (4.3)
Combining (4) and (4), it follows that
$y^{\prime}(t)+c_{1}y(t)+\frac{L_{4}}{4}\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}\leq c_{2}\int_{\Omega}v^{p-r}\hskip 18.49988pt\text{for all }t\in(0,T_{\rm max}).$ (4.4)
Applying Lemma 3.1 with $M=c_{2}+c_{1}$ and $\epsilon=\frac{L_{4}}{4}$, we derive that
$\frac{d}{dt}\int_{\Omega}v^{p-r}+(c_{2}+c_{1})\int_{\Omega}v^{p-r}\leq\frac{L_{4}}{4}\int_{\Omega}\frac{u^{p+1}}{v^{r+1}}+c_{3}\hskip 18.49988pt\text{for all }t\in(0,T_{\rm max}),$ (4.5)
for some $c_{3}>0$. Combining (4.4) and (4.5), we arrive at
$\frac{d}{dt}F_{\lambda}(u,v)+c_{1}F_{\lambda}(u,v)\leq c_{3}\hskip 18.49988pt\text{for all }t\in(0,T_{\rm max}),$
which combines with Gronwall’s inequality implies (4.1) for any $\frac{n}{2}<p<\frac{1+\delta}{\chi^{2}}$. Now, for any $1<q<p$, applying Hölder’s inequality and Lemma 2.2, it follows that
@@
Applying Lemma 4.1 yields there exists $p_{0}\in\left(\frac{n}{2},n\right)$ such that
@@
then we see from Lemma 2.1 that $M(T)<\infty$. Since $p_{0}>\frac{n}{2}$ implies that $\frac{p_{0}+1}{2(n-p_{0}+1)}>\frac{n}{2}$, we can choose $p\in\left(\frac{n}{2},\min\left\{p_{0},\frac{p_{0}+1}{2(n-p_{0}+1)}\right\}\right)$. Employing Hölder’s inequality, we obtain
$\int_{\Omega}u^{p}\leq\left(\int_{\Omega}u^{p_{0}}v^{-r}\right)^{\frac{p}{p_{0}}}\cdot\left(\int_{\Omega}v^{\frac{pr}{p_{0}-p}}\right)^{\frac{p_{0}-p}{p_{0}}}.$
This, together with (4.6), implies that
$\left\|u(\cdot,t)\right\|_{L^{p}(\Omega)}\leq c_{1}^{\frac{1}{p_{0}}}\left\|v(\cdot,t)\right\|_{L^{\frac{pr}{p_{0}-p}}(\Omega)}^{\frac{r}{p_{0}}}\hskip 18.49988pt\text{for all }t\in(0,T_{\rm max}).$ (4.7)